John and Peter play mobile games. Peter posted a high score of 1,450,000 on Temple Run, which is much higher than his previous high score and John suspects Peter is cheating. John has been recording his and Peter's daily scores on the game and calculates that Peter's average score on the game up to this point is 1,100,000 with a standard deviation of 190,000. Assumes Peter's scores are normally distributed.

a) What is the z-score for Peter's new high score? (Round to no less than two decimal places.)

b)Based on his performance up to now, what is the probability that Peter would get that score or higher? (Give the proportion correct to four decimal places.
Answer #1

(a)

= 1100000

= 190000

X = 1450000

Z = (1450000 -1100000)/190000

= 1.8421

(b)

By Technology, Cumulative Area Under Standard Normal Curve = 0.9673

So,

P(X 1450000) = 1 - 0.9673 = 0.0327

So,

Answer is:

0.0327